Question 17 • Algebra 2 with Probability
Since 71% of Earth is water, the probability that one randomly selected location is over land is 1 − 0.71 = 0.29. It is easier to find the probability that all 6 locations are over water: 0.71⁶ ≈ 0.1281. Therefore, the probability that at least 1 location is over land is 1 − 0.1281 = 0.8719, or 0.8719 (choice B).
AI-written and checked against the verified answer.